OpenStudy (jhannybean):

Calculus 3: simple-ish simplification

OpenStudy (jhannybean):

uploading a picture, one sec.

OpenStudy (jhannybean):

I was trying to solve for this but the radical is confusing me.

OpenStudy (anonymous):

so lets for example do d/dy [ z/sqrt(x^2+y^2+z^2) ] this is : z * (-0.5) * 2y * (x^2+y^2+z^2)^(-3/2)

OpenStudy (jhannybean):

Im just unsure how to take the partials. Clarifying my question, haha

OpenStudy (anonymous):

-zy/(x^2+y^2+z^2)^(3/2)

OpenStudy (jhannybean):

Oh I see.

OpenStudy (jhannybean):

Then since all the bases are common, we can factor all of them out.

OpenStudy (anonymous):

what do you mean ? you have to differentiate them

OpenStudy (anonymous):

you mean factor out after the differentiation ?

OpenStudy (jhannybean):

the \((x^2+y^2+z^2)^{3/2}\)

OpenStudy (anonymous):

yes ok

OpenStudy (anonymous):

but you need to differentiate each term with it

OpenStudy (jhannybean):

Yes, i get that :P

OpenStudy (anonymous):

you cant just take it out before the differentiation because there are variables in it

OpenStudy (ankit042):

yup! but wait that is just denominator I think you will get -(yz+xy+xz) in the numerator

OpenStudy (jhannybean):

I was trying to make it simpler for myself instead of including it all within the vector notation, to pull it out as a common factor

OpenStudy (jhannybean):

Ok, so i'll write up what i have, or try to.

OpenStudy (anonymous):

you should get 0 eventually.

OpenStudy (jhannybean):

...I was trying to figure that out on my own, but thanks, lol

OpenStudy (anonymous):

sorry

OpenStudy (jhannybean):

So if all the components derivitate like the one you showed me an example of, can i also factor out the negative from the (-1/2)? I know the 2 is eliminated and al that other stuff.

OpenStudy (anonymous):

yes but remember there are different signs for each differentiation in the curl

OpenStudy (jhannybean):

But we're left with...\[-\frac{1}{(x^2+y^2+z^2)^{3/2}}\] right?

OpenStudy (jhannybean):

the stuff we've factored out i mean!

OpenStudy (anonymous):

you should do all the differentiation first then factor out cause i see that you are getting confused

OpenStudy (jhannybean):

\[\sf -\frac{1}{(x^2+y^2+z^2)^{3/2}}\left< yz - zy~,~-(xz-zx)~,~(xy-yx)\right>\]

OpenStudy (jhannybean):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Then a scalar \(\cdot\) < 0 vector> = \(\hat 0\) right?

OpenStudy (jhannybean):

Or rather, just a 0 vector, lol.

OpenStudy (anonymous):

yes the result of the curl is vector

OpenStudy (jhannybean):

Ack, Im a little confused on how to find the divergence now :( ...

OpenStudy (anonymous):

divergence is easier. you have to take the derivative of the x component with respect to x then add the derivative of the y component with respect to y and then add the same for z

OpenStudy (jhannybean):

Yeah, I understand what you have to do... but it's doing it that's a little... weird I guess.

OpenStudy (anonymous):

what is your first term ?

OpenStudy (jhannybean):

\[\frac{\partial }{\partial x} \left(\frac{x}{(x^2+y^+z^2)^{1/2}}\right)\] of this....

OpenStudy (anonymous):

yes show me what you got

OpenStudy (jhannybean):

Question is... do i simply use the product rule?

OpenStudy (anonymous):

yes and treat x as the only variable

OpenStudy (jhannybean):

\[\frac{\partial}{\partial x} \left[x(x^2+y^2+z^2)^{-1/2}\right]\]\[= 1\cdot (x^2+y^2+z^2)^{-1/2} -\frac{1}{2}(x^2+y^2+z^2)^{-3/2}(2x)\]\[=\frac{x^2+y^2+z^ -x}{(x^2+y^2+z^2)^{3/2}}\]

OpenStudy (jhannybean):

I understood halfway through writing it, heh.

OpenStudy (anonymous):

the bottom line should be y^2+z^2 divided by (x^2+y^2+z^2)^(3/2)

OpenStudy (jhannybean):

I am not sure why the negative etween the z and x seems to be missing, or seems like it, but yeah

OpenStudy (jhannybean):

Bottom line?

OpenStudy (anonymous):

of the d/dx part

OpenStudy (jhannybean):

Oh yes, that's what i was missing, thank you

OpenStudy (anonymous):

so now you know how the d/dy and d/dz will be

OpenStudy (anonymous):

and you know the answer ..?

OpenStudy (jhannybean):

Well analyzing this format, I know that for each component, i will multiply the numerator of the first fraction by \((x^2+y^2+z^2)\), so theonly thing that will change is the variable i'm taking the derivative of.

OpenStudy (jhannybean):

therefore I don't really need to solve for each derivative to understand that I will have:

OpenStudy (jhannybean):

\[=\frac{x^2+y^2+z^2 -x}{(x^2+y^2+z^2)^{3/2}} +\frac{x^2+y^2+z^2 -y}{(x^2+y^2+z^2)^{3/2}}+\frac{x^2+y^2+z^2 -z}{(x^2+y^2+z^2)^{3/2}} \] right?

OpenStudy (anonymous):

no, it is -x^2 instead of -x -y^2 instead of -y -z^2 instead of -z

OpenStudy (jhannybean):

why are they squared now? when I solved for the partial w.r.t x, i got \[=\frac{x^2+y^2+z^2 -x}{(x^2+y^2+z^2)^{3/2}}\]

OpenStudy (jhannybean):

nvm

OpenStudy (anonymous):

but this is wrong i told you you should left with y^2+z^2 divided by (x^2+y^2+z^2)^(3/2)

OpenStudy (jhannybean):

I found my mistake

OpenStudy (anonymous):

good

OpenStudy (jhannybean):

I didn't multiply in the x that was my first function

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you can simplify your answer now

OpenStudy (jhannybean):

\[= 1\cdot (x^2+y^2+z^2)^{-1/2} -\frac{1}{2}(x^2+y^2+z^2)^{-3/2}(2x)(x)\] yeah.

OpenStudy (jhannybean):

And sorry if I'm a bit slow atm, it's 5:30 AM here, heh.

OpenStudy (anonymous):

oh lol go get some sleep

OpenStudy (anonymous):

so what do you get ?

OpenStudy (jhannybean):

\[=\frac{y^2+z^2 }{(x^2+y^2+z^2)^{3/2}} +\frac{x^2+z^2}{(x^2+y^2+z^2)^{3/2}}+\frac{x^2+y^2}{(x^2+y^2+z^2)^{3/2}}\]\[=\frac{2x^2+2y^2+2z^2}{(x^2+y^2+z^2)^{3/2}}\]

OpenStudy (anonymous):

good which is..

OpenStudy (jhannybean):

it's 2 over the square root of the botton function. Haha sorry, you're eager to go. I got it xD

OpenStudy (jhannybean):

Thank you though! You are quite helpful.

OpenStudy (anonymous):

dont want to be rude but i have to leave lol

OpenStudy (anonymous):

so it is correct

OpenStudy (jhannybean):

No i know what you mean, lol. Take care.

OpenStudy (anonymous):

ty you too. have good time

OpenStudy (ikram002p):

still need help ?